LISTSERV mailing list manager LISTSERV 16.0

Help for HEANET-FAULTS Archives


HEANET-FAULTS Archives

HEANET-FAULTS Archives


April 2008


View:

Show Author | Hide Author
Show Table of Contents | Hide Table of Contents

 

Table of Contents:

HEA-NOC/20080306-2 [CLOSE] Maintenance work at TCD Sat 15th March (1 message)
HEA-NOC/20080306-7 [CLOSE] Power maintenance at DKIT on 29th & 30th March 2008. (1 message)
HEA-NOC/20080310-1 [CLOSE] Primary connection to JANET down (1 message)
HEA-NOC/20080310-1 [UPDATE] Primary connection to JANET down (2 messages)
HEA-NOC/20080311-10 [CLOSE] Planned works to affect JANET links (1 message)
HEA-NOC/20080311-10 [UPDATE] Planned works to affect JANET links (1 message)
HEA-NOC/20080313-3 [CLOSE] Network maintenance to DIT Airport building link (1 message)
HEA-NOC/20080313-5 [UPDATE] GÉANT Paris router migration (2 messages)
HEA-NOC/20080314-9 [CLOSE] Power issue in the Galway PoP (1 message)
HEA-NOC/20080320-1 [CLOSE] Mirror server unresponsive (1 message)
HEA-NOC/20080320-1 [UPDATE] Mirror server unresponsive (2 messages)
HEA-NOC/20080320-3 [CLOSE] Migration of ITB to new IP platform (1 message)
HEA-NOC/20080320-3 [UPDATE] Migration of ITB to new IP platform (1 message)
HEA-NOC/20080325-4 [UPDATE] GDA (Grangegorman Development Agency) Migration (2 messages)
HEA-NOC/20080325-5 [UPDATE] ICHEC Migration to New IP equipment (3 messages)
HEA-NOC/20080326-3 [CLOSE] Dublin Core Ring fibre testing (1 message)
HEA-NOC/20080326-3 [UPDATE] Dublin Core Ring fibre testing (3 messages)
HEA-NOC/20080327-5 [CLOSE] NCRI Office Move (1 message)
HEA-NOC/20080327-6 [CLOSE] Loss of resilience on national backbone and to Blanchard. PoP (1 message)
HEA-NOC/20080327-6 [UPDATE] Loss of resilience on national backbone and to Blanchard. PoP (7 messages)
HEA-NOC/20080327-8 [UPDATE] Fault on Videoconferencing ISDN Gateway (1 message)
HEA-NOC/20080328-1 [CLOSE] Connectivity to UL campus lost this morning (1 message)
HEA-NOC/20080330-3 [CLOSE] Grangegorman Outtage (1 message)
HEA-NOC/20080331-2 [CLOSE] Migration of NCRI to New IP Netwrok (1 message)
HEA-NOC/20080331-2 [UPDATE] Migration of NCRI to New IP Netwrok (2 messages)
HEA-NOC/20080331-7 [CLOSE] Etherchannel Bonding on Cork School of Architecture (csa-sw1) (1 message)
HEA-NOC/20080401-10 [CLOSE] Connection to Crumlin VEC down (1 message)
HEA-NOC/20080401-10 [OPEN] Connection to Crumlin VEC down (1 message)
HEA-NOC/20080401-10 [UPDATE] Connection to Crumlin VEC down (6 messages)
HEA-NOC/20080401-4 [OPEN] INEX route server peering (1 message)
HEA-NOC/20080401-4 [UPDATE] INEX route server peering (1 message)
HEA-NOC/20080401-6 [CLOSE] Rom-mon upgrades (1 message)
HEA-NOC/20080401-6 [OPEN] Rom-mon upgrades (1 message)
HEA-NOC/20080401-6 [UPDATE] Rom-mon upgrades (3 messages)
HEA-NOC/20080402-13 [OPEN] Migration to GEANT London Link (1 message)
HEA-NOC/20080402-13 [UPDATE] Migration to GEANT London Link (7 messages)
HEA-NOC/20080402-16 [CLOSE] Scheduled EPA outage (1 message)
HEA-NOC/20080402-16 [OPEN] Scheduled EPA outage (1 message)
HEA-NOC/20080402-16 [UPDATE] Scheduled EPA outage (2 messages)
HEA-NOC/20080403-3 [OPEN] 10G link LAN extension Install (1 message)
HEA-NOC/20080403-3 [UPDATE] 10G link LAN extension Install (2 messages)
HEA-NOC/20080404-11 [CLOSE] Migration of IT Sligo to new IP network (1 message)
HEA-NOC/20080404-11 [OPEN] Migration of IT Sligo to new IP network (1 message)
HEA-NOC/20080404-11 [UPDATE] Migration of IT Sligo to new IP network (2 messages)
HEA-NOC/20080404-8 [OPEN] Scheduled ESBT works affecting LYIT connectivity (1 message)
HEA-NOC/20080404-8 [UPDATE] Scheduled ESBT works affecting LYIT connectivity (6 messages)
HEA-NOC/20080407-3 [OPEN] Degraded service to TippInst, RCSI and NCIRL (1 message)
HEA-NOC/20080407-3 [UPDATE] Degraded service to TippInst, RCSI and NCIRL (4 messages)
HEA-NOC/20080408-1 [CLOSE] Advertise IT Tallaght's Test Range to Smart (1 message)
HEA-NOC/20080408-1 [OPEN] Advertise IT Tallaght's Test Range to Smart (2 messages)
HEA-NOC/20080408-3 [OPEN] Configuring BGP session to cr2-kp for the Services Network (1 message)
HEA-NOC/20080408-3 [UPDATE] Configuring BGP session to cr2-kp for the Services Network (6 messages)
HEA-NOC/20080408-4 [OPEN] Power issues in Blanchardstown Services PoP (1 message)
HEA-NOC/20080408-4 [UPDATE] Power issues in Blanchardstown Services PoP (3 messages)
HEA-NOC/20080408-6 [OPEN] INEX Lan #2 move to new equipment (1 message)
HEA-NOC/20080408-6 [UPDATE] INEX Lan #2 move to new equipment (2 messages)
HEA-NOC/20080409-2 [CLOSE] News Server Errors (1 message)
HEA-NOC/20080409-2 [OPEN] News Server Errors (1 message)
HEA-NOC/20080409-2 [UPDATE] News Server Errors (4 messages)
HEA-NOC/20080409-4 [CLOSE] TCD EduRoam downtime Thursday 10th April (1 message)
HEA-NOC/20080409-4 [OPEN] TCD EduRoam downtime Thursday 10th April (1 message)
HEA-NOC/20080410-1 [OPEN] Loss of connectivity to GTI (1 message)
HEA-NOC/20080410-1 [UPDATE] Loss of connectivity to GTI (3 messages)
HEA-NOC/20080410-12 [OPEN] IP connectivity problems between Mater Dei and DCU campus (1 message)
HEA-NOC/20080410-12 [UPDATE] IP connectivity problems between Mater Dei and DCU campus (1 message)
HEA-NOC/20080410-2 [OPEN] IPv6 connectivity to mirror service disrupted (1 message)
HEA-NOC/20080410-2 [UPDATE] IPv6 connectivity to mirror service disrupted (1 message)
HEA-NOC/20080410-5 [OPEN] NUIG Network Transition Second Phase (1 message)
HEA-NOC/20080410-5 [UPDATE] NUIG Network Transition Second Phase (3 messages)
HEA-NOC/20080411-2 [CLOSE] IUA Router Install (1 message)
HEA-NOC/20080411-2 [OPEN] IUA Router Install (1 message)
HEA-NOC/20080412-1 [CLOSE] Loss of Resiliency- HRB (1 message)
HEA-NOC/20080412-1 [OPEN] Loss of Resiliency- HRB (1 message)
HEA-NOC/20080412-1 [UPDATE] Loss of Resiliency- HRB (1 message)
HEA-NOC/20080412-3 [CLOSE] Loss of Connectivity - Tony Ryan Academy (1 message)
HEA-NOC/20080412-3 [OPEN] Loss of Connectivity - Tony Ryan Academy (1 message)
HEA-NOC/20080412-3 [UPDATE] Loss of Connectivity - Tony Ryan Academy (3 messages)
HEA-NOC/20080413-3 [CLOSE] Testing Between DCU and Materdei (1 message)
HEA-NOC/20080413-3 [OPEN] Testing Between DCU and Materdei (1 message)
HEA-NOC/20080414-8 [CLOSE] updates to keyhole.heanet.ie (1 message)
HEA-NOC/20080414-8 [OPEN] updates to keyhole.heanet.ie (1 message)
HEA-NOC/20080415-2 [CLOSE] CAO Migration to New IP Netwwork (1 message)
HEA-NOC/20080415-2 [OPEN] CAO Migration to New IP Netwwork (1 message)
HEA-NOC/20080416-4 [CLOSE] HEAnet office Loss of resilience (1 message)
HEA-NOC/20080416-4 [OPEN] HEAnet office Loss of resilience (1 message)
HEA-NOC/20080416-7 [OPEN] WIT site Install (1 message)
HEA-NOC/20080416-7 [UPDATE] WIT site Install (1 message)
HEA-NOC/20080417-4 [CLOSE] IUA Migration to New IP Network (1 message)
HEA-NOC/20080417-4 [OPEN] RIA Migration to New IP Network (1 message)
HEA-NOC/20080417-4 [UPDATE] IUA Migration to New IP Network (1 message)
HEA-NOC/20080420-2 [OPEN/CLOSE] Power outage in Confederation House (1 message)
HEA-NOC/20080422-2 [OPEN] New IP transit connection delivery and testing (1 message)
HEA-NOC/20080422-2 [UPDATE] New IP transit connection delivery and testing (3 messages)
HEA-NOC/20080422-3 [CLOSE] EPA Miration to new IP network (1 message)
HEA-NOC/20080422-3 [OPEN] EPA Miration to new IP network (1 message)
HEA-NOC/20080422-3 [UPDATE] EPA Miration to new IP network (1 message)
HEA-NOC/20080422-4 [OPEN] ENET Planned Works Letterkenny IT (1 message)
HEA-NOC/20080422-4 [UPDATE] ENET Planned Works Letterkenny IT (2 messages)
HEA-NOC/20080422-7 [CLOSE] Loss of resilience to NUIM (1 message)
HEA-NOC/20080422-7 [OPEN] Loss of resilience to NUIM (1 message)
HEA-NOC/20080422-7 [UPDATE] Loss of resilience to NUIM (1 message)
HEA-NOC/20080423-2 [CLOSE] IT Tralee Offline (1 message)
HEA-NOC/20080423-2 [OPEN] IT Tralee Offline (1 message)
HEA-NOC/20080423-2 [UPDATE] IT Tralee Offline (2 messages)
HEA-NOC/20080423-5 [CLOSE] HRB physical link down (1 message)
HEA-NOC/20080423-5 [OPEN] HRB physical link down (1 message)
HEA-NOC/20080424-14 [CLOSE] Loss of connectivity to CDVEC Liberties. (1 message)
HEA-NOC/20080424-14 [OPEN] Loss of connectivity to CDVEC Liberties. (1 message)
HEA-NOC/20080424-14 [UPDATE] Loss of connectivity to CDVEC Liberties. (3 messages)
HEA-NOC/20080424-8 [OPEN] Removing UCC Old Links (1 message)
HEA-NOC/20080424-8 [UPDATE] Removing UCC Old Links (1 message)
HEA-NOC/20080424-9 [CLOSE] Eircom - Scheduled work on "unused fiber cable" (1 message)
HEA-NOC/20080424-9 [OPEN] Eircom - Scheduled work on "unused fiber cable" (1 message)
HEA-NOC/20080424-9 [UPDATE] Eircom - Scheduled work on "unused fiber cable" (1 message)
HEA-NOC/20080425-15 [OPEN] UCC cable swap (1 message)
HEA-NOC/20080425-15 [UPDATE] UCC cable swap (4 messages)
HEA-NOC/20080425-16 [OPEN] ESB national fibre ring work (1 message)
HEA-NOC/20080425-6 [CLOSE] Letterkenny IT inaccessible (1 message)
HEA-NOC/20080425-6 [OPEN] Letterkenny IT inaccessible (1 message)
HEA-NOC/20080426-1 [OPEN] Loss of resilience to TippInst (1 message)
HEA-NOC/20080426-1 [UPDATE] Loss of resilience to TippInst (2 messages)
HEA-NOC/20080427-1 [OPEN] Circuits to DIAS Merrion office are down (1 message)
HEA-NOC/20080427-1 [UPDATE] Circuits to DIAS Merrion office are down (1 message)
HEA-NOC/20080428-10 [CLOSE] GEANT Geneva BGP flap (1 message)
HEA-NOC/20080428-10 [OPEN] GEANT Geneva BGP flap (1 message)
HEA-NOC/20080428-10 [UPDATE] GEANT Geneva BGP flap (1 message)
HEA-NOC/20080429-12 [OPEN] GMIT migration to new IP network (1 message)
HEA-NOC/20080429-14 [OPEN] Migrate UL uplink to the new HEAnet core routers (1 message)
HEA-NOC/20080429-14 [UPDATE] Migrate UL uplink to the new HEAnet core routers (2 messages)
HEA-NOC/20080429-2 [OPEN] Janet Backup Scheduled Maintenance (1 message)
HEA-NOC/20080429-8 [OPEN] TCD Outage (1 message)
HEA-NOC/20080429-8 [UPDATE] TCD Outage (1 message)
HEA-NOC/20080430-2 [OPEN] NUIM secondary link brief outage (1 message)
HEA-NOC/20080430-4 [OPEN/CLOSE] Tiscali emergency maintenance, Commodity (1 message)
HEA-NOC/20080430-6 [OPEN] Problems reaching the Inuk IPTV multicast streams (1 message)
HEA-NOC/20080430-6 [UPDATE] Problems reaching the Inuk IPTV multicast streams (1 message)
Test (1 message)

LISTSERV Archives

LISTSERV Archives

HEANET-FAULTS Home

HEANET-FAULTS Home

Subject Sorted by Subject, Chronologically

From

Date

Size

New Thread

HEA-NOC/20080306-2 [CLOSE] Maintenance work at TCD Sat 15th March

HEA-NOC/20080306-2 [CLOSE] Maintenance work at TCD Sat 15th March

Paul Cosgrove

Tue, 1 Apr 2008 14:50:05 +0100

175 lines

New Thread

HEA-NOC/20080306-7 [CLOSE] Power maintenance at DKIT on 29th & 30th March 2008.

HEA-NOC/20080306-7 [CLOSE] Power maintenance at DKIT on 29th & 30th March 2008.

Rob Gallagher

Tue, 1 Apr 2008 10:48:44 +0100

45 lines

New Thread

HEA-NOC/20080310-1 [CLOSE] Primary connection to JANET down

HEA-NOC/20080310-1 [CLOSE] Primary connection to JANET down

Justin Hourigan

Thu, 17 Apr 2008 11:31:39 +0100

71 lines

New Thread

HEA-NOC/20080310-1 [UPDATE] Primary connection to JANET down

HEA-NOC/20080310-1 [UPDATE] Primary connection to JANET down

Justin Hourigan

Mon, 14 Apr 2008 10:17:23 +0100

69 lines

HEA-NOC/20080310-1 [UPDATE] Primary connection to JANET down

Louis Twomey

Fri, 4 Apr 2008 09:15:12 +0100

66 lines

New Thread

HEA-NOC/20080311-10 [CLOSE] Planned works to affect JANET links

HEA-NOC/20080311-10 [CLOSE] Planned works to affect JANET links

Kevin Dermody

Tue, 8 Apr 2008 11:15:14 +0100

39 lines

New Thread

HEA-NOC/20080311-10 [UPDATE] Planned works to affect JANET links

HEA-NOC/20080311-10 [UPDATE] Planned works to affect JANET links

Justin Hourigan

Sun, 6 Apr 2008 12:22:51 +0100

37 lines

New Thread

HEA-NOC/20080313-3 [CLOSE] Network maintenance to DIT Airport building link

HEA-NOC/20080313-3 [CLOSE] Network maintenance to DIT Airport building link

Justin Hourigan

Thu, 17 Apr 2008 11:03:07 +0100

37 lines

New Thread

HEA-NOC/20080313-5 [UPDATE] GÉANT Paris router migration

HEA-NOC/20080313-5 [UPDATE] GÉANT Paris router migration

Dave Wilson

Thu, 10 Apr 2008 17:50:49 +0100

66 lines

HEA-NOC/20080313-5 [UPDATE] GÉANT Paris router migration

Dave Wilson

Thu, 10 Apr 2008 10:14:02 +0100

61 lines

New Thread

HEA-NOC/20080314-9 [CLOSE] Power issue in the Galway PoP

HEA-NOC/20080314-9 [CLOSE] Power issue in the Galway PoP

Justin Hourigan

Mon, 14 Apr 2008 11:41:04 +0100

56 lines

New Thread

HEA-NOC/20080320-1 [CLOSE] Mirror server unresponsive

HEA-NOC/20080320-1 [CLOSE] Mirror server unresponsive

Rob Gallagher

Wed, 9 Apr 2008 17:06:19 +0100

80 lines

New Thread

HEA-NOC/20080320-1 [UPDATE] Mirror server unresponsive

HEA-NOC/20080320-1 [UPDATE] Mirror server unresponsive

Rob Gallagher

Wed, 2 Apr 2008 12:24:02 +0100

77 lines

HEA-NOC/20080320-1 [UPDATE] Mirror server unresponsive

Rob Gallagher

Tue, 1 Apr 2008 10:29:22 +0100

74 lines

New Thread

HEA-NOC/20080320-3 [CLOSE] Migration of ITB to new IP platform

HEA-NOC/20080320-3 [CLOSE] Migration of ITB to new IP platform

Brian Nisbet

Thu, 3 Apr 2008 12:28:01 +0100

48 lines

New Thread

HEA-NOC/20080320-3 [UPDATE] Migration of ITB to new IP platform

HEA-NOC/20080320-3 [UPDATE] Migration of ITB to new IP platform

Brian Nisbet

Wed, 2 Apr 2008 09:39:59 +0100

45 lines

New Thread

HEA-NOC/20080325-4 [UPDATE] GDA (Grangegorman Development Agency) Migration

HEA-NOC/20080325-4 [UPDATE] GDA (Grangegorman Development Agency) Migration

Cormac O'Cianain

Thu, 10 Apr 2008 09:52:58 +0100

42 lines

HEA-NOC/20080325-4 [UPDATE] GDA (Grangegorman Development Agency) Migration

Cormac O'Cianain

Thu, 3 Apr 2008 08:29:17 +0100

40 lines

New Thread

HEA-NOC/20080325-5 [UPDATE] ICHEC Migration to New IP equipment

HEA-NOC/20080325-5 [UPDATE] ICHEC Migration to New IP equipment

Cormac O'Cianain

Wed, 9 Apr 2008 08:55:05 +0100

45 lines

HEA-NOC/20080325-5 [UPDATE] ICHEC Migration to New IP equipment

Cormac O'Cianain

Thu, 3 Apr 2008 08:27:08 +0100

43 lines

HEA-NOC/20080325-5 [UPDATE] ICHEC Migration to New IP equipment

Cormac O'Cianain

Wed, 2 Apr 2008 13:17:14 +0100

37 lines

New Thread

HEA-NOC/20080326-3 [CLOSE] Dublin Core Ring fibre testing

HEA-NOC/20080326-3 [CLOSE] Dublin Core Ring fibre testing

Eoin Kenny

Mon, 28 Apr 2008 16:54:16 +0100

52 lines

New Thread

HEA-NOC/20080326-3 [UPDATE] Dublin Core Ring fibre testing

HEA-NOC/20080326-3 [UPDATE] Dublin Core Ring fibre testing

Brian Boyle

Mon, 28 Apr 2008 15:55:57 +0100

49 lines

HEA-NOC/20080326-3 [UPDATE] Dublin Core Ring fibre testing

Eoin Kenny

Fri, 25 Apr 2008 16:22:01 +0100

47 lines

HEA-NOC/20080326-3 [UPDATE] Dublin Core Ring fibre testing

Eoin Kenny

Fri, 4 Apr 2008 14:18:38 +0100

40 lines

New Thread

HEA-NOC/20080327-5 [CLOSE] NCRI Office Move

HEA-NOC/20080327-5 [CLOSE] NCRI Office Move

Tony Gray

Fri, 4 Apr 2008 11:21:17 +0100

37 lines

New Thread

HEA-NOC/20080327-6 [CLOSE] Loss of resilience on national backbone and to Blanchard. PoP

HEA-NOC/20080327-6 [CLOSE] Loss of resilience on national backbone and to Blanchard. PoP

Tony Gray

Thu, 24 Apr 2008 14:32:02 +0100

72 lines

New Thread

HEA-NOC/20080327-6 [UPDATE] Loss of resilience on national backbone and to Blanchard. PoP

HEA-NOC/20080327-6 [UPDATE] Loss of resilience on national backbone and to Blanchard. PoP

Tony Gray

Tue, 15 Apr 2008 10:25:34 +0100

69 lines

HEA-NOC/20080327-6 [UPDATE] Loss of resilience on national backbone and to Blanchard. PoP

Tony Gray

Mon, 14 Apr 2008 12:22:16 +0100

65 lines

HEA-NOC/20080327-6 [UPDATE] Loss of resilience on national backbone and to Blanchard. PoP

Garwin Liu

Sat, 12 Apr 2008 17:05:49 +0100

62 lines

HEA-NOC/20080327-6 [UPDATE] Loss of resilience on national backbone and to Blanchard. PoP

Garwin Liu

Sat, 12 Apr 2008 14:49:53 +0100

58 lines

HEA-NOC/20080327-6 [UPDATE] Loss of resilience on national backbone and to Blanchard. PoP

Brian Nisbet

Fri, 11 Apr 2008 13:34:42 +0100

52 lines

HEA-NOC/20080327-6 [UPDATE] Loss of resilience on national backbone and to Blanchard. PoP

Tony Gray

Wed, 9 Apr 2008 17:43:10 +0100

45 lines

HEA-NOC/20080327-6 [UPDATE] Loss of resilience on national backbone and to Blanchard. PoP

Tony Gray

Tue, 8 Apr 2008 22:03:23 +0100

42 lines

New Thread

HEA-NOC/20080327-8 [UPDATE] Fault on Videoconferencing ISDN Gateway

HEA-NOC/20080327-8 [UPDATE] Fault on Videoconferencing ISDN Gateway

Brian McArdle

Mon, 31 Mar 2008 16:13:32 +0100

42 lines

New Thread

HEA-NOC/20080328-1 [CLOSE] Connectivity to UL campus lost this morning

HEA-NOC/20080328-1 [CLOSE] Connectivity to UL campus lost this morning

Kevin Dermody

Tue, 8 Apr 2008 15:59:25 +0100

44 lines

New Thread

HEA-NOC/20080330-3 [CLOSE] Grangegorman Outtage

HEA-NOC/20080330-3 [CLOSE] Grangegorman Outtage

Louis Twomey

Tue, 1 Apr 2008 16:18:56 +0100

40 lines

New Thread

HEA-NOC/20080331-2 [CLOSE] Migration of NCRI to New IP Netwrok

HEA-NOC/20080331-2 [CLOSE] Migration of NCRI to New IP Netwrok

Garwin Liu

Wed, 2 Apr 2008 15:49:47 +0100

59 lines

New Thread

HEA-NOC/20080331-2 [UPDATE] Migration of NCRI to New IP Netwrok

HEA-NOC/20080331-2 [UPDATE] Migration of NCRI to New IP Netwrok

Garwin Liu

Wed, 2 Apr 2008 15:49:25 +0100

57 lines

HEA-NOC/20080331-2 [UPDATE] Migration of NCRI to New IP Netwrok

Garwin Liu

Tue, 1 Apr 2008 17:24:21 +0100

55 lines

New Thread

HEA-NOC/20080331-7 [CLOSE] Etherchannel Bonding on Cork School of Architecture (csa-sw1)

HEA-NOC/20080331-7 [CLOSE] Etherchannel Bonding on Cork School of Architecture (csa-sw1)

Garwin Liu

Wed, 2 Apr 2008 08:26:49 +0100

41 lines

New Thread

HEA-NOC/20080401-10 [CLOSE] Connection to Crumlin VEC down

HEA-NOC/20080401-10 [CLOSE] Connection to Crumlin VEC down

Louis Twomey

Thu, 3 Apr 2008 16:35:18 +0100

69 lines

New Thread

HEA-NOC/20080401-10 [OPEN] Connection to Crumlin VEC down

HEA-NOC/20080401-10 [OPEN] Connection to Crumlin VEC down

Gareth Eason

Tue, 1 Apr 2008 20:46:30 +0100

27 lines

New Thread

HEA-NOC/20080401-10 [UPDATE] Connection to Crumlin VEC down

HEA-NOC/20080401-10 [UPDATE] Connection to Crumlin VEC down

Paul Cosgrove

Wed, 2 Apr 2008 19:05:52 +0100

61 lines

HEA-NOC/20080401-10 [UPDATE] Connection to Crumlin VEC down

Louis Twomey

Wed, 2 Apr 2008 17:31:25 +0100

57 lines

HEA-NOC/20080401-10 [UPDATE] Connection to Crumlin VEC down

Louis Twomey

Wed, 2 Apr 2008 11:00:55 +0100

53 lines

HEA-NOC/20080401-10 [UPDATE] Connection to Crumlin VEC down

Louis Twomey

Wed, 2 Apr 2008 10:22:37 +0100

48 lines

HEA-NOC/20080401-10 [UPDATE] Connection to Crumlin VEC down

Gareth Eason

Tue, 1 Apr 2008 21:19:18 +0100

42 lines

HEA-NOC/20080401-10 [UPDATE] Connection to Crumlin VEC down

Gareth Eason

Tue, 1 Apr 2008 21:06:23 +0100

32 lines

New Thread

HEA-NOC/20080401-4 [OPEN] INEX route server peering

HEA-NOC/20080401-4 [OPEN] INEX route server peering

Dave Wilson

Tue, 1 Apr 2008 16:05:43 +0100

33 lines

New Thread

HEA-NOC/20080401-4 [UPDATE] INEX route server peering

HEA-NOC/20080401-4 [UPDATE] INEX route server peering

Dave Wilson

Wed, 2 Apr 2008 10:26:43 +0100

40 lines

New Thread

HEA-NOC/20080401-6 [CLOSE] Rom-mon upgrades

HEA-NOC/20080401-6 [CLOSE] Rom-mon upgrades

Mark Mc Nicholas

Mon, 21 Apr 2008 10:43:01 +0100

61 lines

New Thread

HEA-NOC/20080401-6 [OPEN] Rom-mon upgrades

HEA-NOC/20080401-6 [OPEN] Rom-mon upgrades

Mark Mc Nicholas

Tue, 1 Apr 2008 17:08:54 +0100

45 lines

New Thread

HEA-NOC/20080401-6 [UPDATE] Rom-mon upgrades

HEA-NOC/20080401-6 [UPDATE] Rom-mon upgrades

Mark Mc Nicholas

Wed, 16 Apr 2008 07:56:09 +0100

56 lines

HEA-NOC/20080401-6 [UPDATE] Rom-mon upgrades

Mark Mc Nicholas

Wed, 9 Apr 2008 16:48:49 +0100

53 lines

HEA-NOC/20080401-6 [UPDATE] Rom-mon upgrades

Mark Mc Nicholas

Wed, 9 Apr 2008 16:44:54 +0100

49 lines

New Thread

HEA-NOC/20080402-13 [OPEN] Migration to GEANT London Link

HEA-NOC/20080402-13 [OPEN] Migration to GEANT London Link

Paul Cosgrove

Wed, 2 Apr 2008 17:09:32 +0100

47 lines

New Thread

HEA-NOC/20080402-13 [UPDATE] Migration to GEANT London Link

HEA-NOC/20080402-13 [UPDATE] Migration to GEANT London Link

Dave Wilson

Thu, 10 Apr 2008 12:10:57 +0100

78 lines

HEA-NOC/20080402-13 [UPDATE] Migration to GEANT London Link

Paul Cosgrove

Wed, 9 Apr 2008 09:35:40 +0100

74 lines

HEA-NOC/20080402-13 [UPDATE] Migration to GEANT London Link

Paul Cosgrove

Mon, 7 Apr 2008 16:02:09 +0100

69 lines

HEA-NOC/20080402-13 [UPDATE] Migration to GEANT London Link

Paul Cosgrove

Mon, 7 Apr 2008 11:48:37 +0100

64 lines

HEA-NOC/20080402-13 [UPDATE] Migration to GEANT London Link

Justin Hourigan

Sat, 5 Apr 2008 18:33:43 +0100

57 lines

HEA-NOC/20080402-13 [UPDATE] Migration to GEANT London Link

Paul Cosgrove

Thu, 3 Apr 2008 19:49:20 +0100

55 lines

HEA-NOC/20080402-13 [UPDATE] Migration to GEANT London Link

Paul Cosgrove

Thu, 3 Apr 2008 15:41:22 +0100

52 lines

New Thread

HEA-NOC/20080402-16 [CLOSE] Scheduled EPA outage

HEA-NOC/20080402-16 [CLOSE] Scheduled EPA outage

Louis Twomey

Wed, 9 Apr 2008 13:04:29 +0100

45 lines

New Thread

HEA-NOC/20080402-16 [OPEN] Scheduled EPA outage

HEA-NOC/20080402-16 [OPEN] Scheduled EPA outage

Louis Twomey

Wed, 2 Apr 2008 17:01:13 +0100

33 lines

New Thread

HEA-NOC/20080402-16 [UPDATE] Scheduled EPA outage

HEA-NOC/20080402-16 [UPDATE] Scheduled EPA outage

Louis Twomey

Fri, 4 Apr 2008 16:28:49 +0100

42 lines

HEA-NOC/20080402-16 [UPDATE] Scheduled EPA outage

Louis Twomey

Fri, 4 Apr 2008 13:48:04 +0100

38 lines

New Thread

HEA-NOC/20080403-3 [OPEN] 10G link LAN extension Install

HEA-NOC/20080403-3 [OPEN] 10G link LAN extension Install

Cormac O'Cianain

Thu, 3 Apr 2008 09:46:50 +0100

30 lines

New Thread

HEA-NOC/20080403-3 [UPDATE] 10G link LAN extension Install

HEA-NOC/20080403-3 [UPDATE] 10G link LAN extension Install

Cormac O'Cianain

Fri, 18 Apr 2008 12:00:16 +0100

38 lines

HEA-NOC/20080403-3 [UPDATE] 10G link LAN extension Install

Cormac O'Cianain

Fri, 4 Apr 2008 09:46:36 +0100

35 lines

New Thread

HEA-NOC/20080404-11 [CLOSE] Migration of IT Sligo to new IP network

HEA-NOC/20080404-11 [CLOSE] Migration of IT Sligo to new IP network

Kevin Dermody

Tue, 8 Apr 2008 11:15:41 +0100

45 lines

New Thread

HEA-NOC/20080404-11 [OPEN] Migration of IT Sligo to new IP network

HEA-NOC/20080404-11 [OPEN] Migration of IT Sligo to new IP network

Louis Twomey

Fri, 4 Apr 2008 16:55:11 +0100

33 lines

New Thread

HEA-NOC/20080404-11 [UPDATE] Migration of IT Sligo to new IP network

HEA-NOC/20080404-11 [UPDATE] Migration of IT Sligo to new IP network

NUIM Technical Contacts

Tue, 8 Apr 2008 09:18:37 +0100

47 lines

HEA-NOC/20080404-11 [UPDATE] Migration of IT Sligo to new IP network

Louis Twomey

Tue, 8 Apr 2008 09:18:37 +0100

43 lines

New Thread

HEA-NOC/20080404-8 [OPEN] Scheduled ESBT works affecting LYIT connectivity

HEA-NOC/20080404-8 [OPEN] Scheduled ESBT works affecting LYIT connectivity

Louis Twomey

Fri, 4 Apr 2008 12:43:56 +0100

35 lines

New Thread

HEA-NOC/20080404-8 [UPDATE] Scheduled ESBT works affecting LYIT connectivity

HEA-NOC/20080404-8 [UPDATE] Scheduled ESBT works affecting LYIT connectivity

Daniel Lete

Tue, 29 Apr 2008 14:41:07 +0100

61 lines

HEA-NOC/20080404-8 [UPDATE] Scheduled ESBT works affecting LYIT connectivity

Daniel Lete

Tue, 29 Apr 2008 12:27:12 +0100

56 lines

HEA-NOC/20080404-8 [UPDATE] Scheduled ESBT works affecting LYIT connectivity

Daniel Lete

Mon, 28 Apr 2008 12:15:30 +0100

50 lines

HEA-NOC/20080404-8 [UPDATE] Scheduled ESBT works affecting LYIT connectivity

Garreth Malone

Sun, 27 Apr 2008 11:15:30 +0100

46 lines

HEA-NOC/20080404-8 [UPDATE] Scheduled ESBT works affecting LYIT connectivity

Garreth Malone

Sun, 27 Apr 2008 10:52:08 +0100

44 lines

HEA-NOC/20080404-8 [UPDATE] Scheduled ESBT works affecting LYIT connectivity

Garreth Malone

Fri, 25 Apr 2008 16:13:49 +0100

40 lines

New Thread

HEA-NOC/20080407-3 [OPEN] Degraded service to TippInst, RCSI and NCIRL

HEA-NOC/20080407-3 [OPEN] Degraded service to TippInst, RCSI and NCIRL

Garreth Malone

Mon, 7 Apr 2008 12:28:02 +0100

33 lines

New Thread

HEA-NOC/20080407-3 [UPDATE] Degraded service to TippInst, RCSI and NCIRL

HEA-NOC/20080407-3 [UPDATE] Degraded service to TippInst, RCSI and NCIRL

NUIM Technical Contacts

Mon, 7 Apr 2008 17:54:07 +0100

49 lines

HEA-NOC/20080407-3 [UPDATE] Degraded service to TippInst, RCSI and NCIRL

Garreth Malone

Mon, 7 Apr 2008 17:54:07 +0100

45 lines

HEA-NOC/20080407-3 [UPDATE] Degraded service to TippInst, RCSI and NCIRL

Garreth Malone

Mon, 7 Apr 2008 15:19:25 +0100

41 lines

HEA-NOC/20080407-3 [UPDATE] Degraded service to TippInst, RCSI and NCIRL

Garreth Malone

Mon, 7 Apr 2008 14:50:15 +0100

37 lines

New Thread

HEA-NOC/20080408-1 [CLOSE] Advertise IT Tallaght's Test Range to Smart

HEA-NOC/20080408-1 [CLOSE] Advertise IT Tallaght's Test Range to Smart

Garwin Liu

Wed, 9 Apr 2008 09:20:29 +0100

37 lines

New Thread

HEA-NOC/20080408-1 [OPEN] Advertise IT Tallaght's Test Range to Smart

HEA-NOC/20080408-1 [OPEN] Advertise IT Tallaght's Test Range to Smart

NUIM Technical Contacts

Tue, 8 Apr 2008 09:40:27 +0100

35 lines

HEA-NOC/20080408-1 [OPEN] Advertise IT Tallaght's Test Range to Smart

Garwin Liu

Tue, 8 Apr 2008 09:40:27 +0100

31 lines

New Thread

HEA-NOC/20080408-3 [OPEN] Configuring BGP session to cr2-kp for the Services Network

HEA-NOC/20080408-3 [OPEN] Configuring BGP session to cr2-kp for the Services Network

Orla McGann

Tue, 8 Apr 2008 12:04:48 +0100

34 lines

New Thread

HEA-NOC/20080408-3 [UPDATE] Configuring BGP session to cr2-kp for the Services Network

HEA-NOC/20080408-3 [UPDATE] Configuring BGP session to cr2-kp for the Services Network

Orla McGann

Wed, 9 Apr 2008 17:59:25 +0100

76 lines

HEA-NOC/20080408-3 [UPDATE] Configuring BGP session to cr2-kp for the Services Network

Orla McGann

Tue, 8 Apr 2008 22:05:49 +0100

55 lines

HEA-NOC/20080408-3 [UPDATE] Configuring BGP session to cr2-kp for the Services Network

Orla McGann

Tue, 8 Apr 2008 21:23:19 +0100

51 lines

HEA-NOC/20080408-3 [UPDATE] Configuring BGP session to cr2-kp for the Services Network

Orla McGann

Tue, 8 Apr 2008 21:21:48 +0100

47 lines

HEA-NOC/20080408-3 [UPDATE] Configuring BGP session to cr2-kp for the Services Network

Orla McGann

Tue, 8 Apr 2008 13:24:36 +0100

43 lines

HEA-NOC/20080408-3 [UPDATE] Configuring BGP session to cr2-kp for the Services Network

Orla McGann

Tue, 8 Apr 2008 12:09:45 +0100

40 lines

New Thread

HEA-NOC/20080408-4 [OPEN] Power issues in Blanchardstown Services PoP

HEA-NOC/20080408-4 [OPEN] Power issues in Blanchardstown Services PoP

Garreth Malone

Tue, 8 Apr 2008 13:20:25 +0100

33 lines

New Thread

HEA-NOC/20080408-4 [UPDATE] Power issues in Blanchardstown Services PoP

HEA-NOC/20080408-4 [UPDATE] Power issues in Blanchardstown Services PoP

Kevin Dermody

Tue, 8 Apr 2008 16:46:44 +0100

45 lines

HEA-NOC/20080408-4 [UPDATE] Power issues in Blanchardstown Services PoP

Kevin Dermody

Tue, 8 Apr 2008 14:43:11 +0100

40 lines

HEA-NOC/20080408-4 [UPDATE] Power issues in Blanchardstown Services PoP

Garreth Malone

Tue, 8 Apr 2008 13:28:55 +0100

37 lines

New Thread

HEA-NOC/20080408-6 [OPEN] INEX Lan #2 move to new equipment

HEA-NOC/20080408-6 [OPEN] INEX Lan #2 move to new equipment

Dave Wilson

Tue, 8 Apr 2008 17:31:24 +0100

35 lines

New Thread

HEA-NOC/20080408-6 [UPDATE] INEX Lan #2 move to new equipment

HEA-NOC/20080408-6 [UPDATE] INEX Lan #2 move to new equipment

Dave Wilson

Thu, 10 Apr 2008 09:44:57 +0100

42 lines

HEA-NOC/20080408-6 [UPDATE] INEX Lan #2 move to new equipment

Dave Wilson

Wed, 9 Apr 2008 17:56:50 +0100

39 lines

New Thread

HEA-NOC/20080409-2 [CLOSE] News Server Errors

HEA-NOC/20080409-2 [CLOSE] News Server Errors

Brian Nisbet

Mon, 21 Apr 2008 11:25:00 +0100

48 lines

New Thread

HEA-NOC/20080409-2 [OPEN] News Server Errors

HEA-NOC/20080409-2 [OPEN] News Server Errors

Brian Nisbet

Wed, 9 Apr 2008 11:21:09 +0100

28 lines

New Thread

HEA-NOC/20080409-2 [UPDATE] News Server Errors

HEA-NOC/20080409-2 [UPDATE] News Server Errors

Louis Twomey

Mon, 14 Apr 2008 16:22:26 +0100

44 lines

HEA-NOC/20080409-2 [UPDATE] News Server Errors

Louis Twomey

Mon, 14 Apr 2008 12:55:29 +0100

41 lines

HEA-NOC/20080409-2 [UPDATE] News Server Errors

Brian Nisbet

Wed, 9 Apr 2008 14:31:44 +0100

38 lines

HEA-NOC/20080409-2 [UPDATE] News Server Errors

Brian Nisbet

Wed, 9 Apr 2008 12:24:43 +0100

32 lines

New Thread

HEA-NOC/20080409-4 [CLOSE] TCD EduRoam downtime Thursday 10th April

HEA-NOC/20080409-4 [CLOSE] TCD EduRoam downtime Thursday 10th April

Kevin Dermody

Thu, 10 Apr 2008 09:29:22 +0100

42 lines

New Thread

HEA-NOC/20080409-4 [OPEN] TCD EduRoam downtime Thursday 10th April

HEA-NOC/20080409-4 [OPEN] TCD EduRoam downtime Thursday 10th April

Kevin Dermody

Wed, 9 Apr 2008 15:28:59 +0100

35 lines

New Thread

HEA-NOC/20080410-1 [OPEN] Loss of connectivity to GTI

HEA-NOC/20080410-1 [OPEN] Loss of connectivity to GTI

Garreth Malone

Thu, 10 Apr 2008 09:34:03 +0100

32 lines

New Thread

HEA-NOC/20080410-1 [UPDATE] Loss of connectivity to GTI

HEA-NOC/20080410-1 [UPDATE] Loss of connectivity to GTI

Garreth Malone

Thu, 10 Apr 2008 12:30:01 +0100

43 lines

HEA-NOC/20080410-1 [UPDATE] Loss of connectivity to GTI

Garreth Malone

Thu, 10 Apr 2008 11:47:51 +0100

39 lines

HEA-NOC/20080410-1 [UPDATE] Loss of connectivity to GTI

Garreth Malone

Thu, 10 Apr 2008 09:56:42 +0100

35 lines

New Thread

HEA-NOC/20080410-12 [OPEN] IP connectivity problems between Mater Dei and DCU campus

HEA-NOC/20080410-12 [OPEN] IP connectivity problems between Mater Dei and DCU campus

Garreth Malone

Thu, 10 Apr 2008 18:09:43 +0100

38 lines

New Thread

HEA-NOC/20080410-12 [UPDATE] IP connectivity problems between Mater Dei and DCU campus

HEA-NOC/20080410-12 [UPDATE] IP connectivity problems between Mater Dei and DCU campus

Garreth Malone

Fri, 11 Apr 2008 09:16:44 +0100

43 lines

New Thread

HEA-NOC/20080410-2 [OPEN] IPv6 connectivity to mirror service disrupted

HEA-NOC/20080410-2 [OPEN] IPv6 connectivity to mirror service disrupted

Rob Gallagher

Thu, 10 Apr 2008 09:40:37 +0100

28 lines

New Thread

HEA-NOC/20080410-2 [UPDATE] IPv6 connectivity to mirror service disrupted

HEA-NOC/20080410-2 [UPDATE] IPv6 connectivity to mirror service disrupted

Orla McGann

Thu, 10 Apr 2008 10:59:12 +0100

37 lines

New Thread

HEA-NOC/20080410-5 [OPEN] NUIG Network Transition Second Phase

HEA-NOC/20080410-5 [OPEN] NUIG Network Transition Second Phase

Brian Nisbet

Thu, 10 Apr 2008 11:28:48 +0100

37 lines

New Thread

HEA-NOC/20080410-5 [UPDATE] NUIG Network Transition Second Phase

HEA-NOC/20080410-5 [UPDATE] NUIG Network Transition Second Phase

Brian Nisbet

Mon, 21 Apr 2008 17:02:42 +0100

47 lines

HEA-NOC/20080410-5 [UPDATE] NUIG Network Transition Second Phase

Brian Nisbet

Tue, 15 Apr 2008 17:30:11 +0100

44 lines

HEA-NOC/20080410-5 [UPDATE] NUIG Network Transition Second Phase

Brian Nisbet

Mon, 14 Apr 2008 10:08:09 +0100

41 lines

New Thread

HEA-NOC/20080411-2 [CLOSE] IUA Router Install

HEA-NOC/20080411-2 [CLOSE] IUA Router Install

Garwin Liu

Mon, 14 Apr 2008 11:28:30 +0100

36 lines

New Thread

HEA-NOC/20080411-2 [OPEN] IUA Router Install

HEA-NOC/20080411-2 [OPEN] IUA Router Install

Garwin Liu

Fri, 11 Apr 2008 10:11:56 +0100

29 lines

New Thread

HEA-NOC/20080412-1 [CLOSE] Loss of Resiliency- HRB

HEA-NOC/20080412-1 [CLOSE] Loss of Resiliency- HRB

Garwin Liu

Sun, 13 Apr 2008 18:26:05 +0100

37 lines

New Thread

HEA-NOC/20080412-1 [OPEN] Loss of Resiliency- HRB

HEA-NOC/20080412-1 [OPEN] Loss of Resiliency- HRB

Garwin Liu

Sat, 12 Apr 2008 09:28:26 +0100

31 lines

New Thread

HEA-NOC/20080412-1 [UPDATE] Loss of Resiliency- HRB

HEA-NOC/20080412-1 [UPDATE] Loss of Resiliency- HRB

Garwin Liu

Sat, 12 Apr 2008 17:32:30 +0100

35 lines

New Thread

HEA-NOC/20080412-3 [CLOSE] Loss of Connectivity - Tony Ryan Academy

HEA-NOC/20080412-3 [CLOSE] Loss of Connectivity - Tony Ryan Academy

Garwin Liu

Sun, 13 Apr 2008 18:28:58 +0100

52 lines

New Thread

HEA-NOC/20080412-3 [OPEN] Loss of Connectivity - Tony Ryan Academy

HEA-NOC/20080412-3 [OPEN] Loss of Connectivity - Tony Ryan Academy

Garwin Liu

Sat, 12 Apr 2008 12:02:30 +0100

32 lines

New Thread

HEA-NOC/20080412-3 [UPDATE] Loss of Connectivity - Tony Ryan Academy

HEA-NOC/20080412-3 [UPDATE] Loss of Connectivity - Tony Ryan Academy

Garwin Liu

Sun, 13 Apr 2008 10:28:44 +0100

49 lines

HEA-NOC/20080412-3 [UPDATE] Loss of Connectivity - Tony Ryan Academy

Garwin Liu

Sun, 13 Apr 2008 09:58:14 +0100

43 lines

HEA-NOC/20080412-3 [UPDATE] Loss of Connectivity - Tony Ryan Academy

Garwin Liu

Sat, 12 Apr 2008 12:31:02 +0100

38 lines

New Thread

HEA-NOC/20080413-3 [CLOSE] Testing Between DCU and Materdei

HEA-NOC/20080413-3 [CLOSE] Testing Between DCU and Materdei

Garwin Liu

Sun, 13 Apr 2008 18:30:16 +0100

30 lines

New Thread

HEA-NOC/20080413-3 [OPEN] Testing Between DCU and Materdei

HEA-NOC/20080413-3 [OPEN] Testing Between DCU and Materdei

Garwin Liu

Sun, 13 Apr 2008 10:21:58 +0100

27 lines

New Thread

HEA-NOC/20080414-8 [CLOSE] updates to keyhole.heanet.ie

HEA-NOC/20080414-8 [CLOSE] updates to keyhole.heanet.ie

Aidan Carty

Tue, 15 Apr 2008 11:12:31 +0100

36 lines

New Thread

HEA-NOC/20080414-8 [OPEN] updates to keyhole.heanet.ie

HEA-NOC/20080414-8 [OPEN] updates to keyhole.heanet.ie

Aidan Carty

Mon, 14 Apr 2008 12:07:22 +0100

29 lines

New Thread

HEA-NOC/20080415-2 [CLOSE] CAO Migration to New IP Netwwork

HEA-NOC/20080415-2 [CLOSE] CAO Migration to New IP Netwwork

Garwin Liu

Wed, 16 Apr 2008 09:58:51 +0100

51 lines

New Thread

HEA-NOC/20080415-2 [OPEN] CAO Migration to New IP Netwwork

HEA-NOC/20080415-2 [OPEN] CAO Migration to New IP Netwwork

Garwin Liu

Tue, 15 Apr 2008 09:27:23 +0100

48 lines

New Thread

HEA-NOC/20080416-4 [CLOSE] HEAnet office Loss of resilience

HEA-NOC/20080416-4 [CLOSE] HEAnet office Loss of resilience

Cormac O'Cianain

Wed, 16 Apr 2008 18:06:51 +0100

35 lines

New Thread

HEA-NOC/20080416-4 [OPEN] HEAnet office Loss of resilience

HEA-NOC/20080416-4 [OPEN] HEAnet office Loss of resilience

Cormac O'Cianain

Wed, 16 Apr 2008 14:32:29 +0100

32 lines

New Thread

HEA-NOC/20080416-7 [OPEN] WIT site Install

HEA-NOC/20080416-7 [OPEN] WIT site Install

Cormac O'Cianain

Wed, 16 Apr 2008 16:44:06 +0100

29 lines

New Thread

HEA-NOC/20080416-7 [UPDATE] WIT site Install

HEA-NOC/20080416-7 [UPDATE] WIT site Install

Cormac O'Cianain

Thu, 17 Apr 2008 08:37:58 +0100

32 lines

New Thread

HEA-NOC/20080417-4 [CLOSE] IUA Migration to New IP Network

HEA-NOC/20080417-4 [CLOSE] IUA Migration to New IP Network

Garwin Liu

Wed, 23 Apr 2008 09:08:37 +0100

48 lines

New Thread

HEA-NOC/20080417-4 [OPEN] RIA Migration to New IP Network

HEA-NOC/20080417-4 [OPEN] RIA Migration to New IP Network

Garwin Liu

Thu, 17 Apr 2008 15:42:53 +0100

39 lines

New Thread

HEA-NOC/20080417-4 [UPDATE] IUA Migration to New IP Network

HEA-NOC/20080417-4 [UPDATE] IUA Migration to New IP Network

Garwin Liu

Tue, 22 Apr 2008 08:19:21 +0100

46 lines

New Thread

HEA-NOC/20080420-2 [OPEN/CLOSE] Power outage in Confederation House

HEA-NOC/20080420-2 [OPEN/CLOSE] Power outage in Confederation House

Cormac O'Cianain

Sun, 20 Apr 2008 17:42:53 +0100

28 lines

New Thread

HEA-NOC/20080422-2 [OPEN] New IP transit connection delivery and testing

HEA-NOC/20080422-2 [OPEN] New IP transit connection delivery and testing

Dave Wilson

Tue, 22 Apr 2008 10:00:06 +0100

37 lines

New Thread

HEA-NOC/20080422-2 [UPDATE] New IP transit connection delivery and testing

HEA-NOC/20080422-2 [UPDATE] New IP transit connection delivery and testing

Dave Wilson

Tue, 29 Apr 2008 14:43:29 +0100

51 lines

HEA-NOC/20080422-2 [UPDATE] New IP transit connection delivery and testing

Dave Wilson

Thu, 24 Apr 2008 14:06:42 +0100

47 lines

HEA-NOC/20080422-2 [UPDATE] New IP transit connection delivery and testing

Dave Wilson

Wed, 23 Apr 2008 11:18:03 +0100

43 lines

New Thread

HEA-NOC/20080422-3 [CLOSE] EPA Miration to new IP network

HEA-NOC/20080422-3 [CLOSE] EPA Miration to new IP network

Tony Gray

Thu, 24 Apr 2008 14:31:07 +0100

40 lines

New Thread

HEA-NOC/20080422-3 [OPEN] EPA Miration to new IP network

HEA-NOC/20080422-3 [OPEN] EPA Miration to new IP network

Tony Gray

Tue, 22 Apr 2008 10:32:45 +0100

30 lines

New Thread

HEA-NOC/20080422-3 [UPDATE] EPA Miration to new IP network

HEA-NOC/20080422-3 [UPDATE] EPA Miration to new IP network

Garwin Liu

Wed, 23 Apr 2008 14:22:34 +0100

38 lines

New Thread

HEA-NOC/20080422-4 [OPEN] ENET Planned Works Letterkenny IT

HEA-NOC/20080422-4 [OPEN] ENET Planned Works Letterkenny IT

Garwin Liu

Tue, 22 Apr 2008 11:10:37 +0100

32 lines

New Thread

HEA-NOC/20080422-4 [UPDATE] ENET Planned Works Letterkenny IT

HEA-NOC/20080422-4 [UPDATE] ENET Planned Works Letterkenny IT

Garreth Malone

Sat, 26 Apr 2008 09:18:41 +0100

40 lines

HEA-NOC/20080422-4 [UPDATE] ENET Planned Works Letterkenny IT

Garreth Malone

Sat, 26 Apr 2008 08:26:03 +0100

37 lines

New Thread

HEA-NOC/20080422-7 [CLOSE] Loss of resilience to NUIM

HEA-NOC/20080422-7 [CLOSE] Loss of resilience to NUIM

Garreth Malone

Wed, 23 Apr 2008 17:29:13 +0100

53 lines

New Thread

HEA-NOC/20080422-7 [OPEN] Loss of resilience to NUIM

HEA-NOC/20080422-7 [OPEN] Loss of resilience to NUIM

Garreth Malone

Tue, 22 Apr 2008 15:19:12 +0100

41 lines

New Thread

HEA-NOC/20080422-7 [UPDATE] Loss of resilience to NUIM

HEA-NOC/20080422-7 [UPDATE] Loss of resilience to NUIM

Garreth Malone

Wed, 23 Apr 2008 09:57:42 +0100

46 lines

New Thread

HEA-NOC/20080423-2 [CLOSE] IT Tralee Offline

HEA-NOC/20080423-2 [CLOSE] IT Tralee Offline

Garwin Liu

Fri, 25 Apr 2008 15:58:31 +0100

40 lines

New Thread

HEA-NOC/20080423-2 [OPEN] IT Tralee Offline

HEA-NOC/20080423-2 [OPEN] IT Tralee Offline

Garwin Liu

Wed, 23 Apr 2008 11:11:03 +0100

30 lines

New Thread

HEA-NOC/20080423-2 [UPDATE] IT Tralee Offline

HEA-NOC/20080423-2 [UPDATE] IT Tralee Offline

Garwin Liu

Thu, 24 Apr 2008 09:31:46 +0100

38 lines

HEA-NOC/20080423-2 [UPDATE] IT Tralee Offline

Garwin Liu

Wed, 23 Apr 2008 16:38:04 +0100

34 lines

New Thread

HEA-NOC/20080423-5 [CLOSE] HRB physical link down

HEA-NOC/20080423-5 [CLOSE] HRB physical link down

Cormac O'Cianain

Wed, 23 Apr 2008 14:16:51 +0100

38 lines

New Thread

HEA-NOC/20080423-5 [OPEN] HRB physical link down

HEA-NOC/20080423-5 [OPEN] HRB physical link down

Dave Wilson

Wed, 23 Apr 2008 14:00:13 +0100

33 lines

New Thread

HEA-NOC/20080424-14 [CLOSE] Loss of connectivity to CDVEC Liberties.

HEA-NOC/20080424-14 [CLOSE] Loss of connectivity to CDVEC Liberties.

Garwin Liu

Fri, 25 Apr 2008 15:40:07 +0100

40 lines

New Thread

HEA-NOC/20080424-14 [OPEN] Loss of connectivity to CDVEC Liberties.

HEA-NOC/20080424-14 [OPEN] Loss of connectivity to CDVEC Liberties.

Garwin Liu

Thu, 24 Apr 2008 17:03:48 +0100

28 lines

New Thread

HEA-NOC/20080424-14 [UPDATE] Loss of connectivity to CDVEC Liberties.

HEA-NOC/20080424-14 [UPDATE] Loss of connectivity to CDVEC Liberties.

Garwin Liu

Fri, 25 Apr 2008 09:48:02 +0100

38 lines

HEA-NOC/20080424-14 [UPDATE] Loss of connectivity to CDVEC Liberties.

Garwin Liu

Fri, 25 Apr 2008 09:08:57 +0100

35 lines

HEA-NOC/20080424-14 [UPDATE] Loss of connectivity to CDVEC Liberties.

Garwin Liu

Thu, 24 Apr 2008 17:15:51 +0100

32 lines

New Thread

HEA-NOC/20080424-8 [OPEN] Removing UCC Old Links

HEA-NOC/20080424-8 [OPEN] Removing UCC Old Links

Cormac O'Cianain

Thu, 24 Apr 2008 14:58:01 +0100

29 lines

New Thread

HEA-NOC/20080424-8 [UPDATE] Removing UCC Old Links

HEA-NOC/20080424-8 [UPDATE] Removing UCC Old Links

Cormac O'Cianain

Thu, 24 Apr 2008 16:36:58 +0100

32 lines

New Thread

HEA-NOC/20080424-9 [CLOSE] Eircom - Scheduled work on "unused fiber cable"

HEA-NOC/20080424-9 [CLOSE] Eircom - Scheduled work on "unused fiber cable"

Garwin Liu

Fri, 25 Apr 2008 16:47:15 +0100

43 lines

New Thread

HEA-NOC/20080424-9 [OPEN] Eircom - Scheduled work on "unused fiber cable"

HEA-NOC/20080424-9 [OPEN] Eircom - Scheduled work on "unused fiber cable"

Garwin Liu

Thu, 24 Apr 2008 15:10:15 +0100

37 lines

New Thread

HEA-NOC/20080424-9 [UPDATE] Eircom - Scheduled work on "unused fiber cable"

HEA-NOC/20080424-9 [UPDATE] Eircom - Scheduled work on "unused fiber cable"

Garwin Liu

Thu, 24 Apr 2008 15:35:46 +0100

41 lines

New Thread

HEA-NOC/20080425-15 [OPEN] UCC cable swap

HEA-NOC/20080425-15 [OPEN] UCC cable swap

Dave Wilson

Fri, 25 Apr 2008 16:43:59 +0100

32 lines

New Thread

HEA-NOC/20080425-15 [UPDATE] UCC cable swap

HEA-NOC/20080425-15 [UPDATE] UCC cable swap

Dave Wilson

Mon, 28 Apr 2008 16:45:57 +0100

44 lines

HEA-NOC/20080425-15 [UPDATE] UCC cable swap

Dave Wilson

Mon, 28 Apr 2008 15:50:08 +0100

41 lines

HEA-NOC/20080425-15 [UPDATE] UCC cable swap

Dave Wilson

Mon, 28 Apr 2008 10:07:55 +0100

38 lines

HEA-NOC/20080425-15 [UPDATE] UCC cable swap

Dave Wilson

Mon, 28 Apr 2008 09:40:39 +0100

36 lines

New Thread

HEA-NOC/20080425-16 [OPEN] ESB national fibre ring work

HEA-NOC/20080425-16 [OPEN] ESB national fibre ring work

Dave Wilson

Fri, 25 Apr 2008 16:56:52 +0100

34 lines

New Thread

HEA-NOC/20080425-6 [CLOSE] Letterkenny IT inaccessible

HEA-NOC/20080425-6 [CLOSE] Letterkenny IT inaccessible

Dave Wilson

Mon, 28 Apr 2008 09:45:07 +0100

36 lines

New Thread

HEA-NOC/20080425-6 [OPEN] Letterkenny IT inaccessible

HEA-NOC/20080425-6 [OPEN] Letterkenny IT inaccessible

Dave Wilson

Fri, 25 Apr 2008 10:47:47 +0100

33 lines

New Thread

HEA-NOC/20080426-1 [OPEN] Loss of resilience to TippInst

HEA-NOC/20080426-1 [OPEN] Loss of resilience to TippInst

Garreth Malone

Sat, 26 Apr 2008 08:49:51 +0100

33 lines

New Thread

HEA-NOC/20080426-1 [UPDATE] Loss of resilience to TippInst

HEA-NOC/20080426-1 [UPDATE] Loss of resilience to TippInst

Garreth Malone

Sat, 26 Apr 2008 14:22:13 +0100

44 lines

HEA-NOC/20080426-1 [UPDATE] Loss of resilience to TippInst

Garreth Malone

Sat, 26 Apr 2008 14:10:03 +0100

39 lines

New Thread

HEA-NOC/20080427-1 [OPEN] Circuits to DIAS Merrion office are down

HEA-NOC/20080427-1 [OPEN] Circuits to DIAS Merrion office are down

Brian Boyle

Sun, 27 Apr 2008 22:51:58 +0100

27 lines

New Thread

HEA-NOC/20080427-1 [UPDATE] Circuits to DIAS Merrion office are down

HEA-NOC/20080427-1 [UPDATE] Circuits to DIAS Merrion office are down

Brian Boyle

Sun, 27 Apr 2008 23:24:08 +0100

33 lines

New Thread

HEA-NOC/20080428-10 [CLOSE] GEANT Geneva BGP flap

HEA-NOC/20080428-10 [CLOSE] GEANT Geneva BGP flap

Cillian Sharkey

Tue, 29 Apr 2008 11:08:56 +0100

35 lines

New Thread

HEA-NOC/20080428-10 [OPEN] GEANT Geneva BGP flap

HEA-NOC/20080428-10 [OPEN] GEANT Geneva BGP flap

Cillian Sharkey

Mon, 28 Apr 2008 17:19:01 +0100

28 lines

New Thread

HEA-NOC/20080428-10 [UPDATE] GEANT Geneva BGP flap

HEA-NOC/20080428-10 [UPDATE] GEANT Geneva BGP flap

Cillian Sharkey

Mon, 28 Apr 2008 17:29:33 +0100

32 lines

New Thread

HEA-NOC/20080429-12 [OPEN] GMIT migration to new IP network

HEA-NOC/20080429-12 [OPEN] GMIT migration to new IP network

Tony Gray

Tue, 29 Apr 2008 16:38:09 +0100

46 lines

New Thread

HEA-NOC/20080429-14 [OPEN] Migrate UL uplink to the new HEAnet core routers

HEA-NOC/20080429-14 [OPEN] Migrate UL uplink to the new HEAnet core routers

Paul Cosgrove

Tue, 29 Apr 2008 17:41:21 +0100

45 lines

New Thread

HEA-NOC/20080429-14 [UPDATE] Migrate UL uplink to the new HEAnet core routers

HEA-NOC/20080429-14 [UPDATE] Migrate UL uplink to the new HEAnet core routers

Paul Cosgrove

Wed, 30 Apr 2008 09:52:52 +0100

61 lines

HEA-NOC/20080429-14 [UPDATE] Migrate UL uplink to the new HEAnet core routers

Paul Cosgrove

Wed, 30 Apr 2008 09:24:25 +0100

54 lines

New Thread

HEA-NOC/20080429-2 [OPEN] Janet Backup Scheduled Maintenance

HEA-NOC/20080429-2 [OPEN] Janet Backup Scheduled Maintenance

Cillian Sharkey

Tue, 29 Apr 2008 09:54:47 +0100

28 lines

New Thread

HEA-NOC/20080429-8 [OPEN] TCD Outage

HEA-NOC/20080429-8 [OPEN] TCD Outage

Cillian Sharkey

Tue, 29 Apr 2008 14:18:56 +0100

28 lines

New Thread

HEA-NOC/20080429-8 [UPDATE] TCD Outage

HEA-NOC/20080429-8 [UPDATE] TCD Outage

Cillian Sharkey

Tue, 29 Apr 2008 17:25:20 +0100

33 lines

New Thread

HEA-NOC/20080430-2 [OPEN] NUIM secondary link brief outage

HEA-NOC/20080430-2 [OPEN] NUIM secondary link brief outage

Cillian Sharkey

Wed, 30 Apr 2008 10:54:31 +0100

29 lines

New Thread

HEA-NOC/20080430-4 [OPEN/CLOSE] Tiscali emergency maintenance, Commodity

HEA-NOC/20080430-4 [OPEN/CLOSE] Tiscali emergency maintenance, Commodity

Daniel Lete

Wed, 30 Apr 2008 11:15:19 +0100

27 lines

New Thread

HEA-NOC/20080430-6 [OPEN] Problems reaching the Inuk IPTV multicast streams

HEA-NOC/20080430-6 [OPEN] Problems reaching the Inuk IPTV multicast streams

Orla McGann

Wed, 30 Apr 2008 14:14:39 +0100

29 lines

New Thread

HEA-NOC/20080430-6 [UPDATE] Problems reaching the Inuk IPTV multicast streams

HEA-NOC/20080430-6 [UPDATE] Problems reaching the Inuk IPTV multicast streams

Orla McGann

Wed, 30 Apr 2008 17:50:00 +0100

37 lines

New Thread

Test

Test

Garwin Liu

Thu, 3 Apr 2008 14:49:50 +0100

18 lines

Advanced Options


Options

Log In

Log In

Get Password

Get Password


Search Archives

Search Archives


Subscribe or Unsubscribe

Subscribe or Unsubscribe


Archives

October 2017
September 2017
August 2017
July 2017
June 2017
May 2017
April 2017
March 2017
February 2017
January 2017
December 2016
November 2016
October 2016
September 2016
August 2016
July 2016
June 2016
May 2016
April 2016
February 2016
January 2016
December 2015
November 2015
October 2015
September 2015
August 2015
July 2015
June 2015
May 2015
April 2015
March 2015
February 2015
January 2015
December 2014
November 2014
October 2014
September 2014
August 2014
July 2014
June 2014
May 2014
April 2014
March 2014
February 2014
January 2014
December 2013
November 2013
October 2013
September 2013
August 2013
July 2013
June 2013
May 2013
April 2013
March 2013
February 2013
January 2013
December 2012
November 2012
October 2012
September 2012
August 2012
July 2012
June 2012
May 2012
April 2012
March 2012
February 2012
January 2012
December 2011
November 2011
October 2011
September 2011
August 2011
July 2011
June 2011
May 2011
April 2011
March 2011
February 2011
January 2011
December 2010
November 2010
October 2010
September 2010
August 2010
July 2010
June 2010
May 2010
April 2010
March 2010
February 2010
January 2010
December 2009
November 2009
October 2009
September 2009
August 2009
July 2009
June 2009
May 2009
April 2009
March 2009
February 2009
January 2009
December 2008
November 2008
October 2008
September 2008
August 2008
July 2008
June 2008
May 2008
April 2008
March 2008
February 2008
January 2008
December 2007
November 2007
October 2007
September 2007
August 2007
July 2007
June 2007
May 2007
April 2007
March 2007
February 2007
January 2007
December 2006
November 2006
October 2006
September 2006
August 2006
July 2006
June 2006
May 2006
April 2006
March 2006
February 2006
January 2006
December 2005
November 2005
October 2005
September 2005
August 2005
July 2005
June 2005
May 2005
April 2005
March 2005
February 2005
January 2005
December 2004
November 2004
October 2004
September 2004
August 2004
July 2004
June 2004
May 2004
April 2004
March 2004
February 2004
January 2004
December 2003
November 2003
October 2003
September 2003
August 2003
July 2003
June 2003
May 2003
April 2003
March 2003
February 2003
January 2003
December 2002
November 2002
October 2002
September 2002
August 2002
July 2002
June 2002
May 2002
April 2002
March 2002
February 2002
January 2002
December 2001
November 2001
October 2001
September 2001
August 2001
July 2001
June 2001
May 2001
April 2001
March 2001
February 2001
January 2001
December 2000
November 2000
October 2000
September 2000
August 2000
July 2000
June 2000
May 2000
April 2000
March 2000
February 2000
January 2000
December 1999
November 1999
October 1999
September 1999
August 1999
July 1999
June 1999
May 1999
April 1999
March 1999
February 1999
January 1999
December 1998
November 1998
October 1998
September 1998
August 1998
July 1998
June 1998
May 1998
April 1998
March 1998
February 1998
January 1998
December 1997
November 1997
October 1997
September 1997
August 1997
July 1997
June 1997
May 1997
April 1997
March 1997
February 1997
January 1997
December 1996
November 1996
October 1996
September 1996
August 1996
July 1996
June 1996
May 1996
April 1996
March 1996
February 1996
January 1996
December 1995
November 1995
October 1995
September 1995
August 1995
July 1995
June 1995
May 1995
April 1995
March 1995
February 1995
January 1995
December 1994
November 1994
October 1994
September 1994
August 1994
July 1994
June 1994
May 1994
April 1994
March 1994
February 1994
January 1994
December 1993
November 1993
October 1993
September 1993
August 1993
July 1993
June 1993

ATOM RSS1 RSS2



LISTSERV.HEANET.IE

Secured by F-Secure Anti-Virus CataList Email List Search Powered by the LISTSERV Email List Manager